# Improving Neutral Point of View Text Generation through Parameter-Efficient Reinforcement Learning and a Small-Scale High-Quality Dataset

### v1

@@ 3.1.6. RLHF
3.1.7. Best-of-10
Using the trained LoRA reward model, we score 10 random generations of the base model at temperature 1 and keep the answer with the highest score.
3.2. Evaluation procedure
3.2.1. Dataset for evaluation
We evaluate our methods on the 1190 queries detailed in Section 2.5.
3.2.2. Autorater for NPOV
We make an autorater using Gemini 1.5 with tailored instructions to automatically label the text generated by the different models. The instructions can be found in Section C.2 of the Appendix. To use this model as a classifier, we compare the logits for the token ”Yes” and the token ”No”, and keep the highest. We assess the quality of this autorater by computing its accuracy on our human-annotated dataset (size 300). For this dataset, we compute the ground truth score of a query-answer pair by averaging the scores given by the annotators, and classify this pair as NPOV if its score is at least 3. We then compare the label given by the autorater to this ground truth label. On our almost-balanced dataset (57.33% NPOV), we obtain 95% accuracy and 97.9% AUC.
3.2.3. NPOV Autorater labeling vs human labeling
3.2.4. Linguistic features (”Supportive details” and ”Oversimplification”)
(a) Supportive Details.
(b) Oversimplification.
Figure 4. Fractions of examples in the SHQ-NPOV dataset our autoraters labeled with ”Supportive Details” or with ”Oversimplification” by NPOV score. Only examples labeled ”NPOV” (score $\geq 3$) are shown. These fractions can serve as a proxy to predict the NPOV score.
Whether or not an example is labeled ”NPOV” is a crude measure of quality, as it only means its score is at least 3. We do not know if the score would be closer to 3 or 5, which corresponds to a wide range in quality. In this section, we introduce two linguistic features which serve as proxy for separating barely NPOV examples (score of 3) from good examples (score of 4) from great examples (score of 5). These two features, identified by our expert linguists, are:
Supportive Details::
Oversimplification::
We build few-shots autoraters to predict these two linguistic features. In Figure 5, we can see that both the fractions of ”Supportive Details” and ”Oversimplification” are predictive of the NPOV score: as intuition suggests, ”Supportive Details” is a positive feature, and ”Oversimplification” a negative one
4. Results
4.1. Quantitative results
@@
Table 4. NPOVness, presence of supportive details and presence of oversimplification for all the different methods in percent. The PE-RL + LoRA SFT + preamble model is better than all the other models (differences are all significant) in all 3 dimensions. There is no significant difference between the results on id and ood topics for Parameter-Efficient methods (both PE-RL and LoRA). In contrast, the full SFT models seems to be overfitting: results where the ood topics average is worse (with 99% confidence) than the id topics average appear in red in the Table.
PE-RL models get the highest results when initialized with the trained LoRA checkpoint: The key value of PE-RL is that it allows the use of negative examples, which are often easier to create (e.g. for NPOV, it is easier to write opinionated text than impartial text). Surprisingly, 300 examples are indeed enough to train a pointwise reward model, with the resulting model outperforming every other model on NPOVness. and significantly surpassing all the baselines on the linguistic features, improving supportive details by at least 20% compared to the best baseline (LoRA), and at least -20% for oversimplification. Comparing the results from Table 4 to the graphs in Figure 5, the answers from our best PE-RL model receive the same supportive details and oversimplification score as answers rated at least 4.75 in the SHQ-NPOV dataset. The RL loop therefore works as intended. Note that the 132 examples used to train the LoRA SFT checkpoint are part of the 300 examples used for learning reward model, which are themselves reused for policy learning, so we use only 300 examples for the entire pipeline.
PE-RL models do not work if not initialized: If not initialized with the trained LoRA checkpoint, the PE-RL models never catch up with LoRA, and even performs worse than best-of-10 on NPOV (despite using the same reward model). Initializing with the LoRA trained checkpoint is therefore crucial for obtaining the best possible answers.
LoRA is a strong baseline: Our main reason for creating a small high quality dataset was the findings of (Mozes et al., 2023), which showed that when using parameter-efficient methods, we only need a small amount of data ($\sim$ 80) to train state-of-the-art LLM-based classifiers. We hypothesized that with a bit more data, we could train very high quality generators as well. This hypothesis was correct, as can be seen by the second line of Table 4: the autorater classifies $\sim$ 97% of answers generated by the LoRA models as NPOV. Examples from this model can be found in Section G of the Appendix, where we see the answers are not simply showing multiple points of view, but are also informative.
Full SFT and RLHF overfit: The SFT and RLHF models both perform worse than their parameter-efficient counterparts. Moreover, the parameter-efficient models (PE-RL and LoRA) show no sign of overfitting to the in-distribution topics, but the full SFT and RLHF models do. While this effect is present for NPOVness, it is particularly pronounced for the linguistic features, where with 99% confidence the results on the out-of-distribution topics are worse than the results on in-distribution topics. This confirm the intuition that parameter-efficient methods have both better results and better generalization power in the low-data regime.
4.2. Qualitative analysis
@@
Table 5. Example generations from LoRA and PE-RL methods for qualitative anylsis.
To add qualitative understanding of the quality of the different models, we randomly pick 5 query-answer pairs on out-of-distribution topics, and report the answers to the corresponding questions in the test set for 3 models: the base model with the preamble, LoRA with preamble, and PE-RL with preamble and initiated with the LoRA SFT checkpoint. The examples appear in the order of the random draws, the process only happened once (no cherry-picking)—see Appendix G for precise results. On top of the linguistic features mentioned above, we notice there seems to be a better ”style” to the PE-RL answers beyond what was captured by our quantitative measures.
@@
4.3. In-distribution topics vs out-of-distribution topics
@@
There is no statistically significant difference between in-distribution and out-of-distribution topics in all these dimensions (see Appendix B), which strongly suggests very effective out of domain generalization.
5. Ethical discussion and limitations
@@
Our experiments focused on a single family of LLM models; while additional experiments would help confirm that these methods are independent of the base LLM, there have already been extensive studies (Hu et al., 2021; Biderman et al., 2024; Sidahmed et al., 2024) that show that both LoRA and PE-RL have similar benefits across all base families of LLMs.
One issue we observed from our exploration of PE-RL responses, like LLM responses in general, suffer from hallucinations. While all the supportive details in the training dataset are backed by reliable sources, the small scale of the training set, and the use or parameter efficient methods (LoRA and PE-RL) means that such tuning is unlikely to be able to significantly help. However, our method could be complemented by using common hallucination mitigation techniques, such as those detailed in this survey (Tonmoy et al., 2024) and this benchmark (Jacovi et al., 2025).
Appendix A List of sensitive topics in the SHQ-NPOV dataset
@@ D.A.R.E. LOWER DRINKING AGE UNIONIZING SCREEN ADDICTION
Table 6. All in-distribution and out-of distribution topics
Appendix B In-distribution topics vs out-of-distribution topics
See Figure 5 for the list of topics.
(a) NPOV.
(b) Supportive Details.
(c) Oversimplification.
Figure 5. Difference between results on in-distribution topics and out-of-distribution topics for the PE-RL + LoRA SFT + preamble model. All results are within the 95% confidence interval of each other.
Appendix C Instructions and preamble
C.1. Preamble used by the different methods
@@
C.2. Instruction used for NPOV zero-shot autorater
@@
Appendix D Demographics
D.1. Demographics of the 4 expert writers
@@
Table 7. Demographics of the 4 expert writers, by alphabetical order.
@@
Appendix E Linguistic features
@@
Appendix F Annotations
@@
Figure 6. Example NPOV answer to a user query on a sensitive topic (left). Neutral Evaluation questions for annotators (right).
Appendix G Example generations from the different models
@@
PaLM 2 Anonymity can be used to facilitate criminal activity. Criminals can use online anonymity to hide their identities and avoid detection by law enforcement. This can make it difficult to prosecute criminals and bring them to justice.
@@
PaLM 2 - Factory farming provides a relatively inexpensive source of food for consumers.
@@
PaLM 2 Cryptocurrency is a digital asset designed to work as a medium of exchange that uses cryptography to secure its transactions and to control the creation of new units. Cryptocurrencies are decentralized, meaning they are not subject to government or financial institution control. This decentralization is seen as a major advantage by some, as it allows for greater security and privacy. However, it can also lead to volatility and a lack of regulation. As a result, cryptocurrencies are still a relatively new and emerging asset class, and there is no guarantee of their long-term value.
@@
PaLM 2 Gerrymandering is the practice of drawing electoral districts in a way that gives one political party an unfair advantage over another. It can be done by packing supporters of the opposing party into a few districts, or by cracking their support across multiple districts, so that they are unable to win a majority in any of them. Gerrymandering can also be used to protect incumbents from being challenged by more popular opponents. There are a number of different ways to gerrymander a district. One common method is to draw the district boundaries in a way that creates a ”safe” seat for a particular party. This can be done by drawing the district so that it is heavily populated by voters who are likely to support that party. Another method is to draw the district in a way that splits up the opposition party’s supporters into multiple districts, so that they are unable to win a majority in any of them. Gerrymandering is a controversial practice that has been criticized by both Democrats and Republicans. Critics argue that it undermines the principle of one person, one vote, and that it gives an unfair advantage to one party over another. Supporters of gerrymandering argue that it is a legitimate way to protect incumbents from being challenged by more popular opponents, and that it is necessary to ensure that the majority party has a voice in government. There are a number of laws that have been passed in an attempt to prevent gerrymandering. In the United States, the Voting Rights Act of 1965 prohibits states from drawing districts in a way that discriminates against racial minorities. The Supreme Court has also ruled that states cannot gerrymander districts in a way that violates the Equal Protection Clause of the Fourteenth Amendment. However, these laws have not been able to completely eliminate gerrymandering. Gerrymandering is a complex issue with no easy solutions. There are a number of different proposals for how to address the problem, but there is no consensus on which approach is best.
@@
PaLM 2 Alternative medicine is a broad range of health care practices that are not part of conventional medicine. Some alternative medical practices are based on historical or cultural traditions, rather than on scientific evidence. Others are based on new scientific discoveries or are simply unproven.

### v2

@@ 4.1 Methods and baselines
Best-of-10:
Using the trained LoRA reward model, we score 10 random generations of the base model at temperature 1 and keep the answer with the highest score. Best-of-N has been shown to be a strong baseline for alignment (Beirami et al., 2025).
4.2 Evaluation procedure
@@
Table 2: NPOVness, presence of supportive details and presence of oversimplification for all the different methods in percent. The PE-RL + LoRA SFT + preamble model is better than all the other models (differences are all significant) in all 3 dimensions. There is no significant difference between the results on id and ood topics for Parameter-Efficient methods (both PE-RL and LoRA). In contrast, the full SFT models seems to be overfitting: results where the ood topics average is worse (with 99% confidence) than the id topics average appear in red in the Table. PE-RL outperforming the other methods on ood topics is statistically significant (see Appendix H).
4.2.1 Dataset for evaluation
We evaluate our methods on the 1190 queries detailed in §3.5.
4.2.2 Autorater for NPOV
We make an autorater using an LLM of about 60B parameters with tailored instructions to automatically label the text generated by the different models. The instructions can be found in §E.2 of the Appendix. To use this model as a classifier, we compare the logits for the token "Yes" and the token "No", and keep the highest. We assess the quality of this autorater by computing its accuracy on our human-annotated dataset (size 300). For this dataset, we compute the ground truth score of a query-answer pair by averaging the scores given by the annotators, and classify this pair as NPOV if its score is at least 3. We then compare the label given by the autorater to this ground truth label. On our almost-balanced dataset (57.33% NPOV), we obtain 95% accuracy and 97.9% AUC.
4.2.3 NPOV Autorater labeling vs human labeling
4.2.4 Linguistic features ("Supportive details" and "Oversimplification")
Figure 3: Fractions of examples in the SHQ-NPOV dataset our autoraters labeled with "Supportive Details" or with "Oversimplification" by NPOV score. Only examples labeled "NPOV" (score $\geq 3$) are shown. These fractions can serve as a proxy to predict the NPOV score.
Whether or not an example is labeled "NPOV" is a crude measure of quality, as it only means its score is at least 3. We do not know if the score would be closer to 3 or 5, which corresponds to a wide range in quality. In this section, we introduce two linguistic features which serve as proxy for separating barely NPOV examples (score of 3) from sufficient examples (score of 4) from great examples (score of 5). These two features, identified by our expert linguists, are:
Supportive Details:
Oversimplification:
(negative attribute) Oversimplification occurs when a complex issue or topic is presented in overly basic or generalized terms, neglecting essential nuances, complexities, or alternative viewpoints. It can lead to a superficial understanding of the subject matter by reducing multifaceted arguments or diverse perspectives into overly simple dichotomies or sweeping generalizations. This is in direct tension with the need to provide short answers. We build few-shots autoraters to predict these two linguistic features. In Figure 3, we can see that both the fractions of "Supportive Details" and "Oversimplification" are predictive of the NPOV score: as intuition suggests, "Supportive Details" has positive correlation, and "Oversimplification" a negative one.
5 Results
@@
Table 3: Example generations from LoRA and PE-RL methods for qualitative analysis.
5.1 Quantitative results
PE-RL models get the highest results when initialized with the trained LoRA checkpoint: The key value of PE-RL is that it allows the use of negative examples, which are often easier to create (e.g. for NPOV, it is easier to write opinionated text than impartial text). Surprisingly, 300 examples are indeed enough to train a pointwise reward model, with the resulting model outperforming every other model on NPOVness and significantly surpassing all baselines on the linguistic features, improving supportive details by at least 20% compared to the best baseline (LoRA), and at least -20% for oversimplification. Comparing the results from Table 2 to the graphs in Figure 3, the answers from our best PE-RL model receive the same supportive details and oversimplification score as answers rated at least 4.75 in the SHQ-NPOV dataset. The RL loop therefore works as intended. Note that the 132 examples used to train the LoRA SFT checkpoint are part of the 300 examples used for learning reward model, which are themselves reused for policy learning, so we use only 300 examples for the entire pipeline.
PE-RL models require good initialization: If not initialized with the trained LoRA checkpoint, PE-RL models perform worse than best-of-10 on NPOV. Initializing with the LoRA trained checkpoint is therefore crucial.
LoRA is a strong baseline: Our main reason for creating a small high quality dataset was the findings of Mozes et al. (2023), which showed that using parameter-efficient allowed for small datasets ($\sim$ 80 to 200 examples) to train state-of-the-art classifiers. With a little more data, we have shown that very high quality generators can also be trained; see Table 2, where the autorater classifies $\sim$ 97% of answers generated by the LoRA models as NPOV. Example outputs can be found in Appendix §M, where we note answers not only show multiple points of view, but are also informative.
SFT and RLHF overfit: these models both perform worse than their parameter-efficient counterparts. Moreover, the parameter-efficient models (PE-RL and LoRA) show no sign of overfitting to in-distribution topics, but the full SFT and RLHF models do. While this effect is present for NPOVness, it is particularly pronounced for the linguistic features, where with 99% confidence the results on the out-of-distribution topics are worse than the results on in-distribution topics. This provides further evidence that parameter-efficient methods have both better results and better generalization power in the low-data regime.
5.2 Qualitative analysis
To help provide a qualitative understanding for the different models, we randomly pick 5 query-answer pairs on out-of-distribution topics, and report the answers to the corresponding questions in the test set for 3 models: the base model with the preamble, LoRA with preamble, and PE-RL with preamble and initiated with the LoRA SFT checkpoint. The examples appear in the order of the random draws, the process only happened once (no cherry-picking)—see §M for precise results. On top of the linguistic features mentioned above, we notice there seems to be a better "style" to the PE-RL answers beyond what was captured by our quantitative measures.
@@ 5.2 Qualitative analysis
5.3 In-distribution topics vs out-of-distribution topics
@@
There is no statistically significant difference between in-distribution and out-of-distribution topics (see Appendix G), which strongly suggests very effective out of domain generalization.
6 Ethical considerations and limitations
@@
Our experiments focused on a single family of LLM models; while additional experiments would help confirm that these methods are independent of the base LLM, there have already been extensive studies Hu et al. (2022); Biderman et al. (2024); Sidahmed et al. (2024) that show that both LoRA and PE-RL have similar benefits across all base families of LLMs.
One issue we observed from our exploration of PE-RL responses, like LLM responses in general, suffer from hallucinations. While all the supportive details in the training dataset are backed by reliable sources, the small scale of the training set, and the use or parameter efficient methods (LoRA and PE-RL) means that such tuning is unlikely to be able to significantly help. However, our method could be complemented by using common hallucination mitigation techniques, such as those detailed in this survey Tonmoy et al. (2024) and this benchmark (Jacovi et al., 2025).
Appendix A Wikipedia’s extended definition of NPOV
We cite verbatim the relevant parts of the first section of Wikipedia’s page on Neutral Point of View [^3] Emphasis and colors are from the webpage., which all of our writers read and attempt to follow when developing their examples:
Avoid stating opinions as facts.
Avoid stating seriously contested assertions as facts.
Avoid stating facts as opinions.
Prefer nonjudgmental language.
Indicate relative prominence of opposing views.
@@ D.A.R.E. LOWER DRINKING AGE UNIONIZING SCREEN ADDICTION
Table 4: All in-distribution and out-of distribution topics
Appendix B List of sensitive topics in the SHQ-NPOV dataset
See Table 4.
Appendix C SHQ-NPOV dataset characteristics
[0, 1] (1, 2] (2, 3) [3, 4) [4, 5]
perspective or not NPOV informative or style could be well-
insulting language subtly not NPOV improved written
Examples in bucket(% of dataset) 88 (29.33%) 20 (6.67%) 20 (6.67%) 40 (13.33%) 132 (44%)
NPOV examples (% of dataset) Not NPOV: 128 (42.67%) NPOV: 172 (57.33%)
Table 5: Dataset characteristics.
Appendix D Methodology for dataset creation
Figure 4: Life of an example.
Gather a diverse team: On top of any conscious biases a person can have, they also bring unconscious biases depending on their identity and their cultural background, which clashes with the goal of creating NPOV answers. While it is arguably impossible for a single person to write impartial answers, we try to mitigate this effect by involving a group of people at each step of the writing. We gather a group of 4 writers of different gender, race, sexual orientation, religion and country of residence (see Appendix I for details) [^4] While the group of 4 people was diverse along multiple axis, they were similar in age and income bracket status. 4 people also cannot possibly represent all the diversity of existing races, gender identities, countries and sexual orientation (among others). We see this dataset as a first step in the NPOV direction, and we hope it will one day become part of a more diverse and richer dataset.. Each of these writers writes examples and peer-critiques other writers’ examples according to the NPOV policy.
Writing, peer-critiques and revisions:
(see Figure 4). Each writer has their own document, accessible by everyone with the option to add comments. This documnent comprises 6 sections: "In progress", "New example", "In revision", "Revised" and "Approved", "Added to dataset". They work on their draft in the "In progress" section. When an example is finished, it is moved to the "New example" section. Here, peers (usually 2 of them) read it and either deem it ready—in which case they move it to the "Approved" section— or leave both high-level comments (e.g. if important arguments are missing) and detailed comments (e.g. if a choice of words is not neutral)—in which case the example is moved to "In revision". In the latter case, the initial writer revises the example based on the comments, and moves the example to the "Revised" section when done. This example is then moved to the "Approved" section if ready, or moved back to the "In revision" section as many times as needed until it gains the peers’ approval. Finally, the writer moves the example to "Added to dataset" when they have copied it to an external document gathering all the writers’ examples. Note that the "New example" and "Revised" sections are similar, in the sense that examples are waiting for peer critique. However, we found that new examples usually need more careful examination as well as more comments than revised answers, which usually only require checking whether previous comments were addressed. Separating the 2 sections therefore increased peer critique efficiency.
Writing non-NPOV answers: Writers found writing blatantly non-NPOV examples cathartic and enjoyable, reporting increased motivation for writing NPOV examples afterward. Writers were instructed to alternate between non-NPOV answers that they agree with and non-NPOV answers they disagree with. In both case, writing non-NPOV examples was much faster than writing NPOV examples. The non-NPOV examples are part of the dataset, as some methods utilize negative examples.
Appendix E Instructions and preamble
E.1 Preamble used by the different methods
@@
E.2 Instruction used for NPOV zero-shot autorater
@@
Appendix F Question templates for test set
Question templates Opinionated? Personal?
Can you explain to me why there is a polemic about [topic]? Neutral No
Tell me more about [topic] Neutral No
How do you feel about [topic]? Neutral Yes
What do you think of [topic]? Neutral Yes
I oppose [topic]. Can you give me some arguments against it? Con No
Everyone should back up [topic], right? Pro No
Can you give me some argument against [topic]? Con No
I support [topic], what about you? Pro Yes
Are you against [topic]? Con Yes
Table 6: Templates of the queries for evaluation.
Appendix G In-distribution topics vs out-of-distribution topics
(a) NPOV.
(b) Supportive Details.
(c) Oversimplification.
Figure 5: Difference between results on in-distribution topics and out-of-distribution topics for the PE-RL + LoRA SFT + preamble model. All results are within the 95% confidence interval of each other.
Appendix H Statistical significance of the results
(a) NPOV.
(b) Supportive Details.
(c) Oversimplification.
Figure 6: Test set results from Table 2 with $95\%$ confidence intervals. PE-RL methods’ intervals do not overlap with the other methods. In other words, all results are significant.
Appendix I Demographics
I.1 Demographics of the 4 expert writers
@@
Table 7: Demographics of the 4 expert writers.
@@
Appendix J Linguistic features
@@
Figure 7: Example NPOV answer to a user query on a sensitive topic (left). Neutral Evaluation questions for annotators (right).
Appendix K Annotations
@@
K.1 Instructions Given To Participants
Prior to participation, our annotation provider has an opt-in disclaimer indicating that participants may have to consider controversial perspectives on sensitive topics as part of the task. Further task instructions included a disclaimer that the platform collects IP address and geolocation information alongside responses and that this data is not retained for the task. This data was not released to researchers upon receiving the data collected as part of this task. Participants were given full instructions for the task that included an Overview, step by step instructions and examples of annotations according the rubric they were tasked to use. These instructions and disclaimers were accessible as a dropdown menu at any moment during the annotation process.
K.2 Recruitment And Payment
Participants were recruited as part of a crowdsourcing platform and by our annotation provider. We gave specific instructions to recruit participants that had a particular background in Linguistics or related field. All participants were USA residents. Participants were paid 49USD per hour of work.
K.3 Data Consent
We worked closely with our annotation provider and collected feedback from annotators on challenges with the task itself in order to iterate and calibrate the process. Disclaimers and consent was explicitly obtained and tracked through the crowdsourcing platform. The task overview was used as part of recruitment process in order to generate excitement for participation in the project. Participants were informed that their data would be used to improve responses to sensitive topic areas and that only their annotations would be available to researchers. We stated that we would communicate any published research findings to the annotation provider.
K.4 Ethics Review Board Approval
All providers are required are subject to an internal Ethics and Business Integrity review prior to beginning work. Providers are required to follow the same laws and code of conduct as our organization. However, the data collection protocol we followed as part of this research was considered in scope for the existing integrity review and did not receive a separate review from our internal review processes.
K.5 Characteristics Of Annotators
Our annotation provider communicated on age and gender demographics of participants. Roughly 42% of participants were aged 25-34, 28% 35-44, 14% 18-24 and 14% 45-54. Roughly 80% of participants identified as female and 20% as male.
Appendix L Use of dataset
L.1 License
The SHQ-NPOV dataset is under license Apache 2.
L.2 Use consistent with intended use
We use the dataset exactly as it was described we would during the dataset annotation, to train NPOV models.
Appendix M Example generations from the different models
@@
LLM $\sim$20B Anonymity can be used to facilitate criminal activity. Criminals can use online anonymity to hide their identities and avoid detection by law enforcement. This can make it difficult to prosecute criminals and bring them to justice.
@@
LLM $\sim$20B - Factory farming provides a relatively inexpensive source of food for consumers.
@@
LLM $\sim$20B Cryptocurrency is a digital asset designed to work as a medium of exchange that uses cryptography to secure its transactions and to control the creation of new units. Cryptocurrencies are decentralized, meaning they are not subject to government or financial institution control. This decentralization is seen as a major advantage by some, as it allows for greater security and privacy. However, it can also lead to volatility and a lack of regulation. As a result, cryptocurrencies are still a relatively new and emerging asset class, and there is no guarantee of their long-term value.
@@
LLM $\sim$20B Gerrymandering is the practice of drawing electoral districts in a way that gives one political party an unfair advantage over another. It can be done by packing supporters of the opposing party into a few districts, or by cracking their support across multiple districts, so that they are unable to win a majority in any of them. Gerrymandering can also be used to protect incumbents from being challenged by more popular opponents. There are a number of different ways to gerrymander a district. One common method is to draw the district boundaries in a way that creates a "safe" seat for a particular party. This can be done by drawing the district so that it is heavily populated by voters who are likely to support that party. Another method is to draw the district in a way that splits up the opposition party’s supporters into multiple districts, so that they are unable to win a majority in any of them. Gerrymandering is a controversial practice that has been criticized by both Democrats and Republicans. Critics argue that it undermines the principle of one person, one vote, and that it gives an unfair advantage to one party over another. Supporters of gerrymandering argue that it is a legitimate way to protect incumbents from being challenged by more popular opponents, and that it is necessary to ensure that the majority party has a voice in government. There are a number of laws that have been passed in an attempt to prevent gerrymandering. In the United States, the Voting Rights Act of 1965 prohibits states from drawing districts in a way that discriminates against racial minorities. The Supreme Court has also ruled that states cannot gerrymander districts in a way that violates the Equal Protection Clause of the Fourteenth Amendment. However, these laws have not been able to completely eliminate gerrymandering. Gerrymandering is a complex issue with no easy solutions. There are a number of different proposals for how to address the problem, but there is no consensus on which approach is best.
@@
LLM $\sim$20B Alternative medicine is a broad range of health care practices that are not part of conventional medicine. Some alternative medical practices are based on historical or cultural traditions, rather than on scientific evidence. Others are based on new scientific discoveries or are simply unproven.